4 big dictionaries, each 10.0 cm thick and each 1.50 kg, are resting side by side flat on a table 1.06 m high.

How much work would it take to stack them one atop the other?

How high would that raise the center-of-mass of all books?


1
Expert's answer
2021-11-04T10:17:43-0400

"m = 1.5kg"

"d = 10cm = 0.1m"

"h_0 = 1.06m"

"g = 9.8\frac{m}{s^2}"

"\text{Move 2 dictionary to 1 dictionary:}"

"W_1 = mg*d"

"\text{Move 3 dictionary to a stack of 1 and 2 dictionaries}"

"W_2 = mg*2d"

"\text{\text{Move 4 dictionary to a stack of 1 and 2 and dictionaries}}"

"W_2 = mg*3d"

"W = W_1+W_2+W_3= mg*6d=1.5*9.8*6*0.1=8.82J"

"W = 4mg\Delta h_c"

"4m - \text{The total mass of dictionaries}"

"\Delta h_c - \text{changing the position of the center of gravity of dictionaries}"

"\Delta h_c=\frac{W}{4mg}=\frac{8.82}{4*9.8*1.5}=0.15m"


"\text{Answer:}W = 8.82J;\Delta h_c = 0.15m"